OpenStudy (anonymous):

What are the roots of x^3 -5x^2 -25x +125 and what kind are they?

OpenStudy (luigi0210):

You could solve this by grouping: \[\LARGE x^3-5x^2=x^2(x-5)\] and \[\LARGE -25x+125=-25(x-5)\] Now just combine: \[\LARGE (x^2-25)(x-5)\] Factor again: \[\LARGE (x-5)(x+5)(x-5)\] Combine and get: \[\LARGE (x-5)^2(x+5)\]

OpenStudy (anonymous):

Both of them are real, and not complex, correct?

OpenStudy (luigi0210):

Yup, they're real solutions.

OpenStudy (anonymous):

Thank you, very helpful explanation